# SPIN-ODE: Stiff Physics-Informed Neural ODE for Chemical Reaction Rate Estimation

### v2

@@ IV-D Robustness studies
TABLE III: Robustness of CRNN pre-training under increasingly sparse observations. We downsample the data samples from the trajectories on the AOXID dataset, and report the MAE loss on ODE fitting and rate coefficients estimation in SPIN-ODE-s2.
V Conclusion
We propose a novel neural network-based framework for estimating chemical rate coefficients without the need for prior knowledge. To the best of our knowledge, this is the first approach capable of addressing stiff chemical ODEs under these conditions. This capability is particularly important in real-world chemical systems, such as atmospheric chemistry, where rate coefficients can vary exponentially and significantly influence reaction pathways. Our method employs a three-stage optimisation process: it begins by fitting stiff ODE trajectories, proceeds with explicit learning of the complete chemical system using a Chemical Reaction Neural Network (CRNN), and concludes with a refinement step that integrates CRNN with an ODE solver to enhance prediction accuracy. We validate our approach on both synthetic datasets and simplified real-world atmospheric chemical scenarios. The experimental results, both quantitative and qualitative, demonstrate that our method achieves state-of-the-art performance in ODE trajectory fitting and rate coefficient estimation. In future work, we aim to extend the approach to comprehensive chemical models and improve its robustness under more extreme stiffness conditions.
Acknowledgment
The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. W. P. thanks the Onni ja Hilja Tuovinen Foundation Grant for their generous support.

### v3

@@ 4.4 Robustness studies
Table 3: Robustness of CRNN pre-training under increasingly sparse observations. We downsample the AOXID trajectories and report the ODE fitting and rate coefficient estimation errors in SPIN-ODE-s2.
5 Conclusion
We propose a neural network framework for estimating chemical reaction rate coefficients without relying on prior knowledge. To the best of our knowledge, this is the first approach to address stiff chemical ODEs under such conditions. This capability is particularly relevant for real-world systems such as atmospheric chemistry, where rate coefficients vary exponentially and strongly influence reaction rates. Our method follows a three-stage optimisation process: it begins with trajectory fitting of stiff ODEs, proceeds with explicit learning of the full chemical system using CRNN, and concludes with a refinement step that integrates the CRNN with an ODE solver to improve accuracy. We validate the framework on both synthetic datasets and simplified atmospheric scenarios, with results demonstrating state-of-the-art performance in trajectory fitting and rate coefficient estimation. Future work will extend the method to comprehensive chemistry schemes and further strengthen its robustness under real-world conditions.
{ack}
The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. W. P. gratefully acknowledges financial support from the Finnish Cultural Foundation (Päijät-Häme Regional Fund 70251819) and the Onni ja Hilja Tuovinen Foundation Grant.